OpenStudy (anonymous):

L^-1{1/s^2 - 48/s^5}

OpenStudy (anonymous):

please try to make your question clear .tq

OpenStudy (apples):

\[\mathcal{L}^{-1}\{\frac{1}{s^2} - \frac{48}{s^5}\} = \mathcal{L}^{-1}\{\frac{1}{s^2}\} - \mathcal{L}^{-1}\{\frac{48}{s^5}\} = t - 48 \left[\mathcal{L}^{-1}\{\frac{1}{s^5}\}\right]\] which is \[t - 2t^4\]